A baseball player hits a ball such that as it leaves the bat, it has speed of
40 meters per second and has an angle of 440
. What is the time that the ball
reaches the highest point of its flight?
1
Expert's answer
2021-08-27T08:26:04-0400
"t=\frac{v\sin{44}}{g}\\t=\frac{40\sin{44}}{9.8}=2.8\ s"